The shorter leg of a right triangle is 6m shorter than the longer leg. The hypotenuse is 6m longer than the longer leg. Find the side lengths of the triangle.
Find the length of shorter leg, find the length of longer leg and the length of hypotenuse

Found 2 solutions by jorel1380, ikleyn:
Answer by jorel1380(3719)

Let the short leg be n. Then the longer leg would be n+6; and the hypotenuse would be n+12. Then:
n²+(n+6)²=(n+12)²
n²+n²+12n+36=n²+24n+144
n²-12n-108=0
(n-18)(n+6)=0
n=18 or -6
☺☺☺☺

Answer by ikleyn(52780)

.
(3-4-5) right-angled triangle 18 m, 24 m and 30 m.